OpenStudy (anonymous):

Im having trouble with this problem. Write a slope-intercept equation for a line with a slople 1/5 which passed through the point (10,-4). I believe the answer is y=1/5x-4.. but im not sure

OpenStudy (angela210793):

y+4=1/5(x-10) y=x/5-2-4 y=x/5-6